OpenStudy (anonymous):

i thinkof a no & add 12 then i start again with the no double it & subtract 8 .the two results are same. can someone help

OpenStudy (anonymous):

add eight over and subtract x \[4=x\]

Parth (parthkohli):

Another way:\[x - 4 = 2(x - 4)\]So obviously, \(x - 4 = 0\).

OpenStudy (shubhamsrg):

Or, x= +inf or -inf! ;)

Parth (parthkohli):

I hope any one of you didn't think of dividing both sides by \(x - 4\), or you could have been jailed . . .

Parth (parthkohli):

Oh, new question?\[x + 12 = 2x - 8\]

OpenStudy (anonymous):

thanks A lot to help me but please can u tell me how to solve it step by step i ll highly apperciate thanx & regards in advance

OpenStudy (anonymous):

So you want to isolate x. First step is to get like terms on the same side with each other. SO add 8 over and subtract so you get \[4=x\] in this case. It won't always isolate quite so easily there may be division/multiplication involved but the process involved is always the same. Combine like terms and then isolate x

OpenStudy (anonymous):

thanks & regards